### v2

# Unveiling Interpretability in Self-Supervised Speech Representations for Parkinson’s Diagnosis

Abstract
Recent works in pathological speech analysis have increasingly relied on powerful self-supervised speech representations, leading to promising results. However, the complex, black-box nature of these embeddings and the limited research on their interpretability significantly restrict their adoption for clinical diagnosis. To address this gap, we propose a novel, interpretable framework specifically designed to support Parkinson’s Disease (PD) diagnosis. Through the design of simple yet effective cross-attention mechanisms for both embedding- and temporal-level analysis, the proposed framework offers interpretability from two distinct but complementary perspectives. Experimental findings across five well-established speech benchmarks for PD detection demonstrate the framework’s capability to identify meaningful speech patterns within self-supervised representations for a wide range of assessment tasks. Fine-grained temporal analyses further underscore its potential to enhance the interpretability of deep-learning pathological speech models, paving the way for the development of more transparent, trustworthy, and clinically applicable computer-assisted diagnosis systems in this domain. Moreover, in terms of classification accuracy, our method achieves results competitive with state-of-the-art approaches, while also demonstrating robustness in cross-lingual scenarios when applied to spontaneous speech production.
Index Terms:
Parkinson’s Disease, Deep Learning, Cross-Attention Mechanisms, Interpretability, Self-Supervised Speech.I Introduction
Parkinson’s Disease (PD) is the second most common neurodegenerative disorder after Alzheimer’s Disease [nussbaum2003disease]. Furthermore, its prevalence has been observed to increase substantially over the past two decades, underscoring the growing impact of this chronic condition in our society [zhu2024temporal]. The disease is characterized by a progressive loss of neurons, especially in regions of the brain responsible for dopamine production [hornykiewicz1998biochemical, aarsland2017cognitive]. Among the primary consequences of this dopamine deficiency is a decline in motor control, often manifesting in symptoms such as muscular rigidity, tremors, and slowness of movement, alongside a variety of non-motor issues, predominantly related to cognitive, sleep, and sensory abnormalities [jankovic2008parkinson, poewe2017parkinson].
It is estimated that 70% to 90% of individuals with PD develop vocal and speech impairments [logemann1978frequency, hartelius1994speech, miller2007prevalence], including hoarse, breathy voice quality, reduced variation in pitch and loudness, weakened stress, imprecise slurred articulation, and rapid bursts of speech interrupted by inappropriate periods of silence [canter1963intensity, canter1965articulation, darley1969clusters]. Because speech production involves complex coordination between cognitive and physiological processes, these impairments represent valuable indicators that support an automated, cost-effective, and non-invasive diagnosis of PD, among other motor speech disorders [duffy2012motorspeech].
Numerous studies in pathological speech literature have demonstrated evidence of phonatory impairments [hanson1984cinegraphic, zwirner1992vocal, hammer2010laryngeal], articulatory difficulties [canter1965articulation, logemann1981vocal, ackermann1991articulatory], and prosodic deficiencies [canter1963intensity, skodda2008rhythm, jones2009prosody] in individuals with PD, and how these factors affect their communication with others and quality of life [ho1999speech, schalling2018speech, miller2006life]. Indeed, these broader, high-level concepts provide a practical and interpretable starting point for assessing speech impairments, enabling clinicians to subsequently address and refine their focus on specific dimensions of speech pathology as required for diagnosis, screening, and treatment.
Unfortunately, traditional assessments in clinical settings can be highly subjective and time-consuming [ginanneschi1988evaluation, richards1994interrater, ramig2008speechtreatment, vasquez2018towards], highlighting the need for more objective and automated methods for evaluating speech impairments. In this sense, a growing body of research [morovelazquez2021review, rusz2024prodromal] proposes machine learning techniques for analyzing pathological speech patterns, aiming to improve both the efficiency and accuracy of these assessments. While a wide range of approaches has been explored in the field, recent works have shifted the focus from classical, clinically informed speech features [vasquez2020disvoice, eyben2016gemaps] to pre-trained self-supervised learned (SSL) speech representations [babu2202wav2vec, shor2022trillsson, hsu2021hubert]. These SSL methods have shown notable advancements across various speech analysis tasks [bayerl2022stuttering, wagner2023pathological, baumann2024cleftlip], offering promising potential to enhance the diagnostic process of PD [favaro2023interpretable, laquatra2024exploiting, gallo2024levels, escobar2024foundation].
However, despite the improved performance, SSL-based speech representations present notable limitations in interpretability. Due to their complex, often opaque structure, understanding how these models capture specific speech characteristics — such as phonatory, articulatory, or prosodic features — remains particularly challenging. In clinical settings, this black-box nature of SSL models raises important concerns, as diagnostic support systems must be interpretable and explainable to promote trust, transparency, and compliance with ethical and legal standards [rasheed2022explainable]. While recent studies have started to explore interpretability in SSL models broadly [pasad2021layerwise, pasad2023comparative, singla2022whatdo], research specifically tailored to pathological speech is still quite limited [baumann2024cleftlip, escobar2024foundation, gallo2024levels]. In this line, although there is an ongoing debate regarding the reliability of attention mechanisms for interpretability [jain2019not, wiegreffe2019notnot, bibal2022debate], attention-based model interpretations have recently shown promising achievements in studies on cleft lip and palate pathological speech [ariasvergara2023measuring, baumann2024cleftlip, baumann24cleftattn]. Indeed, these methods not only provided interpretable insights, but also present a more resource-efficient alternative to their gradient-based counterparts, making them especially appealing for practical use in clinical settings.
All the aforementioned considerations motivated our research, which aims to address the critical gaps in the interpretability of SSL-based speech models for pathological speech analysis, particularly in the context of Parkinson’s Disease. The key contributions are:
We propose a novel interpretable framework based on cross-attention mechanisms for PD diagnosis support111https://github.com/david-gimeno/interpreting-ssl-parkinson-speech. This framework not only identifies meaningful speech patterns within SSL-based speech representations, but also achieves results competitive with the state-of-the-art.
By addressing model interpretability from two distinct perspectives, our framework uncovers first insights into the speech information encoded by these black-box representations when discerning pathological speech conditions. Besides, it provides fine-grained analyses of how different speech dimensions evolve over time, with a focus on clinical relevance to guide diagnostic procedures.
We validate the effectiveness of our proposed interpretable method and its transparency in model decision-making through extensive experimentation on five well-established benchmarks for PD detection, encompassing a diverse range of speech assessment tasks and languages.
We extend our analysis to cross-language studies, demonstrating promising results in spontaneous speech production, particularly in monologue tasks. This finding underscores the robustness of our attention-based approach and its potential for advancing multilingual PD assessment.
II Related Work
This section provides a brief overview of the current state of the art in PD speech analysis and discusses interpretable deep learning approaches in the field of pathological speech.
Automatic PD Detection from Speech. Speech analysis for PD has been extensively studied over the past decades, resulting in a vast body of research [morovelazquez2021review, rusz2024prodromal]. Despite the wide range of explored approaches, recent works in the field have increasingly adopted powerful self-supervised speech representations, achieving promising results. Favaro et al. [favaro2023interpretable] presented a thorough comparative study between the use of informed, knowledge-based features and SSL-based speech representations. Their findings concluded that, while both types of features exhibited robustness in cross-lingual scenarios, SSL-based representations consistently achieved superior detection accuracies. La Quatra et al. [laquatra2024exploiting] took a further step by exploiting the potential of these foundational speech models and investigating the effects of speech enhancement techniques under real-world operating conditions, where acoustic factors may not be ideal. The diverse range of machine learning methods and deep SSL-based speech models studied make these works particularly representative of the current state of the art in the field, yielding an average F1-score of around 80% in their best-performing and most optimistic settings. However, the evaluation over specific subsets of speech assessment tasks or the varying task grouping strategies adopted in the literature hinders comparisons between methods. Hence, in our work, we propose a task-specific assessment framework. This approach not only facilitates direct comparisons with our method, but is also better suited for interpretability in clinical settings.
Interpretability in Pathological Speech Analysis. A primary objective in this area has been to support the diagnosis of pathological speech conditions through the design of interpretable models [morovelazquez2021review]. While earlier approaches successfully relied on more classical, informed features [orozco2018neurospeech], research on the interpretability of SSL-based speech representations remains limited. Here we discuss studies focusing on explaining deep representations from various perspectives. For instance, Abderrazek et al. [abderrazek2023neuroconcept] proposed a method based on phoneme-level neuron activation patterns in deep convolutional networks to assess intelligibility and phonetic alterations in patients with head and neck cancer. Other studies, similar to ours, utilized frame-level Wav2Vec latent representations to identify phonatory impairments at various linguistic granularity levels, as well as to examine the discriminative potential of different phonological classes (e.g., fricatives, nasals, plosives) for PD [escobar2024foundation, gallo2024levels]. However, unlike our work, these methods not only require phoneme-aligned annotated data for training, but also focus primarily on phonatory features, thus neglecting other clinically relevant speech dimensions. Baumann et al. [baumann2024cleftlip, baumann24cleftattn], in turn, explored, among other gradient-based techniques, the design of interpretable methods based on attention mechanisms in the context of cleft lip and palate speech. While their promising findings inspired our present work, their method still does not consider other relevant speech features, such as prosody or glottal articulation. The modeling of these characteristics poses several challenges when the interpretability mainly depends on phoneme-related information. The key difference and contribution of our proposed framework thus stems from the use of cross-attention mechanisms to integrate the knowledge encoded by informed speech features with the potential of SSL-based representations, thereby extending the applicability of this type of interpretable architectures.
III Method
This section outlines the background and key concepts behind our approach, followed by a detailed description of the interpretable deep learning model designed to detect PD from voice and speech patterns. The overall framework and its underlying motivations are illustrated in Figure 1.
III-A Background
Attention-Based Mechanisms. Attention-based mechanisms, introduced by Vaswani et al. [vaswani2017attention], have demonstrated their powerful capabilities across a variety of tasks and domains, such as natural language processing, computer vision, speech processing, and healthcare, [transformers2024comprehensive]. By allowing the model to capture long-range dependencies and focus on the most relevant parts within the input sequential data, attention mechanisms enhance the model’s ability to detect subtle patterns, making them particularly useful for speech processing tasks such as the one addressed in this study. Mathematically, the attention mechanism can be described as follows:
where , , and refer to the query, key, and value vectors, respectively, and is a normalization factor which corresponds to the dimensionality of the keys. These vectors are derived through linear projections of the input feature embeddings using learnable weights: , , and . Once the input embeddings have been projected, the operation enables each time step in the sequence to attend to all other time steps, thereby allowing the model to capture long-range dependencies. The resulting attention scores are then applied to the value vectors, focusing on the most relevant parts of what, in our study, would be the input speech sequence.
Depending then on how the input to the attention-based module is setup, we can distinguish between:
Self-Attention. In this case, the same input sequence is used to compute the three entries of the attention module, thus capturing relationships within the input itself. This approach is employed for our baseline models, when only one type of speech representation is considered.
Cross-Attention. In contrast, when the input for the queries differs from the keys, the resulting attention scores represent the relationships between two distinct types of representations. These scores then complement and enrich the value vectors, , which correspond to one of the representations used for the queries or keys. This mechanism is employed for our proposed interpretable framework, when both the SSL-based speech representation and informed speech dimensions are considered.
Although a multi-headed attention variant exists, which divides the computation of the attention scores across multiple attention heads, we opted to use the simpler yet effective single-head attention mechanism described above. The rationale behind this choice is that having more independent attention heads could obscure and complicate our goal of interpretability, as each attention head may focus on different aspects of the input sequences. Baumann et al. [baumann24cleftattn] exemplifies the challenges of using multi-head attention in pathological speech analysis. Their findings indicate that each attention head tend to model speech patterns in diverse ways, occasionally resulting in overlapping attention focus, that some heads consistently showed low attention weights, and that the ablation of these heads could even improve model performance — observations that are misaligned with the interpretability and diagnostic support objectives of the present work.
III-B Interpretability through Cross-Attention Mechanisms
Overview. Given a dataset of samples, , where denotes the input speech feature representations and their corresponding pathological label, the goal is to find the optimal parameters for a deep learning model, , by minimizing the average cross entropy loss, , over the entire dataset. Formally, this optimization process can be expressed as: , which is directly tied to the detection of PD at the utterance level.
Interpretability Goals. Beyond model accuracy, our primary motivation was to design a novel framework based on cross-attention mechanisms capable of injecting interpretability into SSL-based speech embeddings, which, besides being powerful latent representations, often function as black boxes.
To achieve this, our proposed model uses both SSL-based speech embedding sequences and informed speech feature sets as inputs. Specifically, each utterance sample is represented by an SSL embedding sequence, , which captures temporal speech patterns in the data, along with an informed feature set, , which contains more interpretable but static speech characteristics. Here, denotes the number of time steps, is the embedding dimension, and represents the number of clinically informed speech features.
Our goal is therefore to obtain two attention score matrices that reflect the relationships between both speech representations. The first attention matrix, , would address interpretability at the embedding level, where for each dimension of the SSL-based embedding, we quantify the contribution of each of the knowledge-based, informed features. Similarly, we aim to obtain a second attention matrix, , which would offer a complementary temporal-level perspective, providing the contribution of each of the informed features at each time step of the SSL sequence.
Model Architecture. Having outlined the interpretability goals of our model, we now turn to the technical details of its architecture. As described above and illustrated in Figure 1, both speech representations are used to compute two types of cross-attention scores within separate and dedicated modules. The input to each attention-based module is defined as follows:
where denote learnable linear projections, while corresponds to the identity matrix , representing a non-transformative operation that leaves the informed features unchanged. This choice, along with the decision to preserve the dimensionality of the SSL-based representations intact, was made intentionally to support our goal of interpretability within the proposed framework. Therefore, in general terms, the clinically informed speech features act directly as the key values, providing a set of interpretable dimensions against which SSL-based embeddings are aligned. However, our dual-branch architecture was designed to investigate model interpretability from two distinct, yet complementary, perspectives. Here is where the static nature of our informed speech features plays a crucial role, offering flexibility to expand them alongside either the temporal or embedding feature dimensions. Depending on the interpretability perspective addressed, different aspects, including slight modifications of the original attention mechanism — specifically matrix translations to align shapes for the derived dot products — were considered:
Embedding Cross-Attention. This module focuses on understanding the information embedded by SSL-based speech representations and their internal structure. To inject interpretability into these otherwise opaque, black-box SSL embeddings, we first repeated the informed speech features times, such that . We then compute the scaled dot-product attention, so all this knowledge is integrated across the temporal dimension. The resulting attention scores, , reveal the overall relationships of each dimension of the SSL-based embedding with each of the informed, more interpretable speech features. These scores are finally applied to the vector to obtain the enriched representation . Ultimately, this module aims to address a critical research gap regarding the type and extent of information encoded within pre-trained, SSL-based representations in the context of pathological speech analysis.
Temporal Cross-Attention. This module, instead, focuses on a fine-grained interpretation of the speech signal, providing insights at each time step. In this case, we first repeated the informed speech features times, such that . We then compute the scaled dot-product attention, integrating all this knowledge across the embedding dimension. The resulting attention scores, , reveal the overall relationships of each temporal step of the SSL-based embedding sequence and the informed, more interpretable speech features. These scores are finally applied to the vector to obtain the enriched representation . Ultimately, this module aims to provide valuable support for clinicians in diagnosing speech pathologies, such as PD.
Once we obtain the enriched, conditioned SSL-based speech representation both at embedding () and temporal () levels, we can proceed with the final classification. First, each conditioned latent sequence is reduced by averaging along its temporal and embedding dimension, respectively. The resulting embeddings are then concatenated, yielding a single embedding . This utterance-level representation is then processed by the classification module, which consists of a linear layer, preceded by layer normalization [ba2016layer] and Swish activation [swish2017prajit], that projects data into a two-dimensional space to perform the target binary classification task.
Discussion. The reliability of attention-based methods for interpretability remains an active area of research [jain2019not, wiegreffe2019notnot, bibal2022debate]. Recent studies in pathological speech analysis, however, have demonstrated promising results, showing alignment between attention-based attributions and clinical expert decisions in cleft lip and palate speech [baumann2024cleftlip, baumann24cleftattn]. These studies, like our work, leverage attention weights to identify potential articulatory and phonatory impairments, supporting the validity of attention mechanisms as a means to identify clinically relevant speech characteristics. Despite these promising insights, the interpretation of these methods entails several considerations. As our framework is trained using an objective loss function directly tied to the final classification of PD, the learned attention scores in an individual sample likely embed information relevant to the presence (or absence) of the pathology, as we later discuss in Section LABEL:sec:embedding. However, since these are attention scores, we cannot guarantee that a high (or low) attention for a specific speech dimension—such as the average duration of pauses—directly corresponds to higher (or lower) occurrence of that feature. Rather, attention scores indicate where the model is allocating more (or less) focus, providing a degree of transparency that clinicians can leverage to guide and support their diagnostic decisions.
Another aspect to discuss is the inclusion or exclusion of linear projections within the cross-attention modules. Specifically, we chose to omit the linear layer for the key values, , in Eq. (2), given that applying this layer poses a risk of obfuscating or altering the structured, informed speech feature set — a concern directly tied to our goal of enhancing model interpretability. Conversely, we retained the linear projection for the query and value inputs, and , respectively, to benefit from their adaptive flexibility. Since the reduced existing research on the internal structure of the SSL-based speech representation remains limited to layer-wise analysis [pasad2023comparative, singla2022whatdo], we cannot assume that speech features are necessarily embedded as isolated values, but they may instead emerge from intricate, latent relationships. Thus, maintaining these projections not only aligns with established attention-based mechanisms [vaswani2017attention], but also allows the model to adaptively learn and uncover relationships within the SSL-based embeddings.
IV Feature Extraction
This section outlines the two speech feature representations employed in our research study. We first introduce the black-blox, SSL-based speech embeddings to subsequently describe the set of more interpretable speech features, discussing and motivating their relevance within our proposed framework.
IV-A Self-Supervised Speech Features
In this work, we employed the 24-layer, 300M-parameter XLR-S Wav2Vec2.0 model [babu2202wav2vec] to extract our SSL-based speech representations222https://pytorch.org/audio/main/generated/torchaudio.pipelines.WAV2VEC2_XLSR_300M.html. Specifically, we extracted 1024-dimensional embeddings at a sample rate of 20ms from the 7th encoder layer. Notably, we used the pre-trained model without fine-tuning for any downstream tasks. Therefore, we are considering a SSL model trained on 436k hours of unlabeled typical-speech audio data from multiple datasets covering 128 languages, including the ones considered in this study.
Although other architectures have also been explored [favaro2023interpretable, laquatra2024exploiting], the choice of this foundational speech model was motivated by its wide adoption in pathological speech analysis. This model, or its variants, were used to assess different speech affecting diseases, including PD [favaro2023interpretable], cleft lip and palate [baumann2024cleftlip], speech impairments resulting from oral or laryngeal cancer [wagner2023pathological], and detecting dysfluencies in stuttering [bayerl2022stuttering]. The success and effectiveness of these latent speech representations across multiple domains reflects that this model is particularly well-suited to the task addressed in our work, providing robust, well-established support for interpretability studies.
Furthermore, by preserving the pre-trained, general speech foundation model without any type of fine-tuning, we exploit its capacity to detect deviations and markers of speech impairments with respect to a healthy population—a crucial aspect in our present intepretability study. Moreover, its multilingual foundation enhances its suitability for our research, allowing us to assess how well such SSL models generalize when applied to pathological speech analysis across diverse languages.
Finally, existing literature on layer-wise analysis of SSL-based foundational speech models [pasad2021layerwise] suggests that these architectures exhibit an autoencoder-like behavior, with deeper layers becoming progressively more abstract, while the final layers start to resemble the input speech features, almost as if reconstructing them. These same studies also found that shallow, intermediate layers tend to capture more acoustic, phonetic, and linguistic information. Consistent with these findings, recent works in the context of pathological speech analysis [favaro2023interpretable, purohit2025layeradapt] have demonstrated that intermediate layer representations can achieve performance comparable to fine-tuning the entire encoder architecture, thus providing a more parameter-cost efficient alternative. Our preliminary layer-wise analysis also revealed that using the 7th layer consistently outperformed using other layers, by a larger margin when compared to extreme layers (e.g., the first and last), and by a smaller margin to adjacent layers. In conclusion, all these findings further support our approach as an effective representation for interpretability in the context of PD diagnosis.
IV-B Interpretable Clinically Informed Speech Features
As introduced throughout the paper, the primary objective of our research work consists of injecting interpretability into black-box, SSL speech features through cross-attention mechanisms. Therefore, the interpretability of our proposed deep-learning model architecture depends entirely on the interpretable, clinically informed speech dimensions we consider.
Our first experiments utilized the full range of articulation, glottal, phonation, and prosody features provided by the DisVoice toolkit [vasquez2020disvoice]. However, this initial feature set comprising 655 features was considered excessively broad for practical interpretability, as the sheer volume of information would be challenging and overwhelming. Moreover, several of these features are not clinically meaningful, as they do not directly relate to anatomical or physiological mechanisms of speech production, thereby limiting their utility for clinical explainability. One example would be the Mel Frequency Cepstral Coefficients, which, despite attempts in some studies [illner2023aspects, tracey2023towards] to interpret them through correlations with more interpretable features, remain inherently non-clinically interpretable.
To provide a more straightforward yet comprehensive identification of these pathological patterns within our proposed method, we selected a reduced set of 35 informed speech dimensions, focusing on those that are clinically meaningful or that are supported by studies that have demonstrated their effectiveness as speech biomarkers of PD. This selection emphasizes features that not only encapsulate a general speech characteristic but also provide a finer level of interpretability, enhancing their practical value for clinicians. Interestingly, reducing the feature set achieved comparable results without significant drops in model performance and, under certain conditions, even improved classification accuracy. Our selected features, thoroughly described in [vasquez2020disvoice], can be broadly categorized into the following four high-level speech characteristics: